Josh stored an electric potential energy of 2.5j in a vacuum of 1.00m3 by a field E. If the field is reduced to 0.25 of its initial value, how much energy is stored per cubic meter?


1
Expert's answer
2021-05-23T18:47:12-0400

we know that ,

electric potential energy in a volume is given by:


(energy density) "\times" (volume) = "({1\over2} \epsilon_0 E^2)\times(volume)"

here,

P.E.=2.5 J= "({1\over2} \epsilon_0 E^2)\times(1)"


now we reduced E by 0.25 times ,

now new potential energy will be,

P.E.= "({1\over2} \epsilon_0({ E\over4})^2)\times(1)"


="2.5\over16" J

=0.156 J